Hey I need help with this tricky question.. thanks in advance!
For a Simple Harmonic Motion, d=(3/4)sin8πt
find:
(a) frequency
(b) the least positive value of t for which d = 0

Answer by Alan3354(69443)

For a Simple Harmonic Motion, d=(3/4)sin8πt
find:
(a) frequency
Period = 2pi/(8pi) = 1/4 seconds/cycle
f = 1/p = 4 cycles/sec
= 4 Hz
=============
(b) the least positive value of t for which d = 0
d(t) = 0 at t = 0
d(t) = 0 at t = 0.125 seconds